Chez Calvin sold 692 pairs of their designer jeans last year at $250.99 each. This year their price increased to $304.99 but units sold were down to 651. Answer the following questions, indicating a decrease with a negative percent change. Note: Please make sure your final answer(s) are in percentage form and are accurate to 2 decimal places. For example 34.56%

What is the percent change in price year-over-year?

Percent Change = 0.00%


What is the percent change in units sold year-over-year?

Percent Change = 0.00%


What is the percent change in total revenue year-over-year?
Answer #1

Percent change in price year over year = difference in prices / old prices * 100

= 54 / 250.99 * 100

= 21.51 %

Percent change in sales year over year = difference in sales / previous year sales * 100

= - 41 / 692 * 100

=( - 5.92 %)

Revenue in previous year = 692*250.99 = 173685.08

Revenue in current year = 651 * 304.99 = 198548.49

Change in percent of revenue = increase in revenue / last years revenue * 100

= 24863.41 / 173685.08 * 100

= 14.31 %